OpenStudy (anonymous):

How many grams of magnesium chloride is produced when .500 moles of magnesium reacts with an excess of hydrochloric acid?

OpenStudy (aaronq):

First write and balance an equation for the reaction. Then use the stoichiometric coefficients to find moles produced. Set up a ratio using the species of interest, like so: e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients , \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) From here you can isolate what you need. For example: if you have 2 moles of B, how many moles of C can you produce? solve algebraically: \(\dfrac{2}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\rightarrow n_C=\dfrac{2*\color{green}{c}}{\color{blue}{b}}\) -------------------------------------------------------- To interconvert from mass to moles or moles to mass, use the relationship: \(n=\dfrac{m}{M}\) where, M=molar mass, m=mass, and n= moles. --------------------------------------------------

OpenStudy (anonymous):

o-o

OpenStudy (anonymous):

Well you could just do \[n =\frac{ M }{ Mr }\] That would give you N x Mr = M 0.500 x 94.3 (Molecular mass of MgCl2) = 47.15 Grams

OpenStudy (aaronq):

^while you're right it doesn't really show how you arrived at that answer, because you can only do that because the stoichiometric ratios are 1:1. If this person assumes that every question is solved like that, then they're not really learning correct the process.

OpenStudy (anonymous):

Of course ratios play a massive part in this but im just assuming that this reaction is all 1:1 like you said. Its good you pointed that out to them, I should have been a better chemist and mentioned that.